three masses are hung on a force table. they are 60 grams hung at 20 degrees, 20 grams hung at 88 degrees, and 45 grams hung at 152 degrees. find where a forth mass must be hung and how much the mass has to be to. balance the ring in the center of the force table


1
Expert's answer
2021-03-31T07:14:09-0400

Let's first find the "x"- and "y"-components of the resultant force:


"\sum F_x=F_{1x}+F_{2x}+F_{3x},""60\ g\cdot cos20^{\circ}+20\ g\cdot cos88^{\circ}+45\ g\cdot cos152^{\circ}=17.35\ g,""\sum F_y=F_{1y}+F_{2y}+F_{3y},""60\ g\cdot sin20^{\circ}+20\ g\cdot sin88^{\circ}+45\ g\cdot sin152^{\circ}=61.63\ g."

We can find the magnitude of the resultant force from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2}=\sqrt{(17.35\ g)^2+(61.63\ g)^2}=64\ g."

We can find the direction from the geometry:


"\theta=cos^{-1}(\dfrac{F_x}{F})=cos^{-1}(\dfrac{17.35\ g}{64\ g})=74.3^{\circ}."


To find the angle on the force table we must subtruct "74.3^{\circ}" from "360^{\circ}":


"\theta=360^{\circ}-74.3^{\circ}=285.7^{\circ}."

Therefore, the 64 g mass at 285.7 degrees will balance the ring in the center of the force table.